You are working on a wood project in woodshop. You have a wooden rod that is 72 inches long. You need to cut the rod so that one piece is 6 inches longer than the other pieces. How long should each piece be ???
Answer by benazir.sj@gmail.com(109)

total length of the wooden rod =72in
length of first rod =x in
length of second rod =(x+6)in
x+(x+6) =72
2x+6 =72
2x=72-6
2x= 66
x=33
x+6 =39
length of the rods are 33in and 39 in